INAF LBC+LUCIFER queue observing, 2011 June 30 -July 01

Observer: R. Speziali, K. Boutsia
Telescope Operator: S. Allanson
Support Astronomer: O. Kuhn
Instrument: LBC

Summary

Clear skies at the beginning of the night.
We intended to start with Lucifer but shortly before sunset we found out that lucifer was warming up because the compressor did not start up after the power outage.
During sunset we switched to LBC. At the beginning of the night it took almost 1hr to collimate the telescope.
We had to repeat our first OB (lost ~45min) due to rotator problems. Apparently the CMU was 2 min late and this was creating problems presetting the rotator path. After the update everything worked fine. In the meantime we changed the exp time of the OBs to 300sec instead of 600sec in order to get less elongated stars.
Later we get the 600 sec exposure back again after we saw round stars on the 300 sec images.

Data Summary

Project Objects Filter ExpTime Notes
Lanzoni Pal14_F1 gSLOAN - rSLOAN 1hr15min exp time changed to 300sec because of rotator problems which caused elongation
Lanzoni Pal14_F2 gSLOAN - rSLOAN 1hr15min exp time back to the original request - rotator problem seems fixed - target completed
Lanzoni GC6205_F1 gSLOAN - rSLOAN 25min central part fo the globular cluster all saturated
Lanzoni GC6205_F2 gSLOAN - rSLOAN 10min stopped after 10 min
Lanzoni GC6779 gSLOAN - rSLOAN 75min seeing within specs - target completed
Lanzoni GC6779 gSLOAN - rSLOAN 15min just for fun - preparing flat OB
Flats

Details

02:31 - Temperature problems with Lucifer
  • the compressor did not start up after the power outage
  • it might take some hours to cool down
  • we decide to switch to LBC

02:41 - Switch to LBC

02:49 - Turn ON LBC
  • we will try for flats - if not good enough use the ones from UT 20110627 or 20110624
  • taking some bias frames - 10Bias_Bino

03:00 - slew for gr_pa0 flats
  • gsloan - 1sec / rsloan 2sec - PA_0
  • gsloan - 1.5sec / rsloan 3sec - PA_180

03:13 - Slew to ACT0247 for collimation and pointing correction
  • the temperature dropped 10deg and the mirrors are not in equilibrium...
  • it seems that we are hitting focus limits...
  • dofpia, /First - we are intrafocal again on the blue side
  • we interrupted - gave some global offsets manually
  • dofpia, /First
  • we manually give Z11=-700 on the red side
  • we exit after the 11th iteration

03:41 - we are roughly collimated and we go for rangebalance
  • there is a 2deg difference in temperature between mirrors and the ambient.

03:46 - Slew to Pal14_F1_2
  • dofpia
  • DIMM ~1.10''
  • exit after the 8th iteration
  • dofopia, - we stop after the 8th iteration - red still not completely converged

04:13 - Pal14_F1_1 - observe
  • we decide to repeat the whole programm because the data from UT 20110627 are out of specs
  • after the first 600sec we see the imaged elongated!! seeing is <1.0''
  • it is a rotational elongation - the rotator is not moving at the correct rate...

04:27 - the second image is elongated as well
  • we turn OFF the camera and On again
  • according to the issueTrak last time this fixed the problem

04:38 - Turn ON LBC - connect to LBT

04:45 - we go on with the 3rd exposure of the previous OB to see if the situation has changed
  • this image looks fine and the sseing is within the specs

04:59 - repeat Pal14_F1_1
  • so that we obtain the first 2 images
  • the last blue seems elongated - the red seems fine

05:22 - dofpia
  • it collimated after 2 iterations

05:25 - Pal14_F1_2
  • we changed exposure time to 300secx6 because we have rotator problems and this way the eventual trail will be smaller
  • DIMM ~1.0''
  • still images on the blue side seem elongated - especially on the 3rd chip

06:00 - dofpia

06:03 - interrupt dofpia to adjust the click on the windows computers that are delaying by 2 min (phone call to N.Cushing)

06:11 - we resume dofpia
  • the lbc control computer times are changed but not on the windows computers
  • we stopped it after the 5th iteration

(note> from now on everything will run smooth!)

06:18 - Pal14_F1_3
  • the exp times are all changed to 300sec

06:36 - dofpia
  • 2 iterations to converge

06:39 - Pal14_F2_1
  • here we kept the original exptime 600sec because the rotator problem should have been solved...
  • the first image is not elongated.
  • seeing measured on the image ~0.9''

07:12 - dofpia
  • 2 iterations to converge

07:16 - Pal14_F2_2 * DIMM seeing 1.1''

07:50 - dofpia
  • 3 iterations to converge

07:54 - Pal14_F2_3 * DIMM seeing 1.1-1.2''

8:11 - Preset to 6205 F1

08:13 - dofpia
  • 4 iterations to converge

08:20 -GC6205_F1_1
  • DIMM seeing 1.1''

08:48 - dofpia
  • stopped after 3 iterations

08:53 - GC6205_F2_1
  • OB acquisition stopped after the 2nd exposure. Next Target has priority.

09:06 - Preset to NGG6779

09:08 - dofpia
  • stopped after 4 iterations (Blue converged after 2)

09:15 - GC6779_1
  • IMG seeing 0.8''

09:44 - dofpia
  • Converged after 3 iterations (Blue converged after 2)

09:50 - GC6779_2
  • IMG seeing 0.8''

10:19 - GC6779_3
  • IMG seeing 0.8''

10:49 - dofpia
  • Converged after 3 iterations

10:55 - GC6779_3 bis
  • IMG seeing 0.8''
  • we stopped during the 4th iteration because the blue channel was completely saturated

11:19 - slewing for flats
  • g-SLOAN - 5sec / r-SLOAN - 10sec - PA_0
  • g-SLOAN - 4sec / r-SLOAN - 8sec - PA_180
  • g-SLOAN - 2sec / r-SLOAN - 4sec - PA_180
  • g-SLOAN - 0.8sec / r-SLOAN - 1.6sec - PA_0 - changed to 0.5sec- 1.0sec because counts too high

11:50 - End of Night

-- KonstantinaBoutsia - 01 Jul 2011
Topic revision: r5 - 01 Jul 2011, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback